OpenStudy (aaronandyson):

If 5 cos (theta) = 3 Evaluate cosec(theta) - cot(theta)/cosec(theta) + cot(theta) @perl

OpenStudy (aaronandyson):

@paki

OpenStudy (perl):

$$ \large 5 \csc (\theta ) = 3 \\ \large \csc(\theta) = \frac {3}{5} $$

OpenStudy (aaronandyson):

Okay, now ?

OpenStudy (irishboy123):

this is a 3,4,5 pythagorean triple. simplest -- just draw the triangle and read off the ratios.

OpenStudy (aaronandyson):

Umm? @perl what did he say?

OpenStudy (perl):

he said good things :) $$ \Large \text { is this question}\\ \Large \frac{(\csc \theta - \cot \theta)}{(\csc \theta + \cot \theta)} $$

OpenStudy (perl):

does it say what the angle of theta is ,

OpenStudy (aaronandyson):

No.

OpenStudy (perl):



OpenStudy (aaronandyson):

missing side is 4? now what

OpenStudy (aaronandyson):

now?

OpenStudy (perl):

$$ \Large \cos \theta = 3/5\\ \Large \sec \theta = 3/4\\ \Large \sin \theta = 4/5 \\ \Large \csc \theta = 5/4 \\ \Large \tan \theta = 4/3\\ \Large \cot \theta = 3/4\\ $$

OpenStudy (aaronandyson):

I mean how to I substitute the values in the expression?

OpenStudy (perl):

$$ \text {Plug into }~ \frac{(\csc \theta - \cot \theta)}{(\csc \theta + \cot \theta)} \\ \huge \frac{\frac{5}{4}-\frac{3}{4}}{\frac{5}{4}+\frac{3}{4}} $$

OpenStudy (aaronandyson):

1?

OpenStudy (aaronandyson):

2/4/8/4

OpenStudy (perl):

$$ \huge \frac{\frac{5}{4}-\frac{3}{4}}{\frac{5}{4}+\frac{3}{4}} = \frac{\frac{2}{4}}{\frac{8}{4}} = \frac{2}{8}=\frac{1}{4} $$

OpenStudy (aaronandyson):

I got 1/4 also :)

OpenStudy (aaronandyson):

help me at more trig sum??

OpenStudy (perl):

ok

OpenStudy (aaronandyson):

@perl

OpenStudy (aaronandyson):

I posted a new question and tagged you.